OpenStudy (anonymous):

Find the length of the radius of the following circle. (x + 2)2 + y2 = 10

OpenStudy (anonymous):

Circle equation is given by: \[(x + h)^2 + (y-k)^2 = r^2\] where r is the radius.. So here : \[r^2 = 10\] So what will be the radius?

OpenStudy (anonymous):

I HATE THIS STUFFFFFFFFFFFFFF LOL!!!

OpenStudy (anonymous):

(10)?

OpenStudy (anonymous):

*(x-h)^2

OpenStudy (anonymous):

or even √(10)

OpenStudy (anonymous):

Yep, this looks right..